Solve the first-order differential equation = \frac{dy}{dx} = \frac{x^{2}+xy+y^{2}}{x^2} with y(-1)=1
Answer #1

\frac{dy}{dx}\:=\:\frac{\left(x^2+xy+y^2\right)}{x^2}
\mathrm{Substitute\quad }\frac{dy}{dx}\mathrm{\:with\:}y'\:
y'\:=\frac{\left(x^2+xy+y^2\right)}{x^2}

\mathrm{Substitute}\:y=ux,\:u\:\mathrm{is\:a\:function\:of}\:x
\left(ux\right)'\:=\frac{x^2+xux+\left(ux\right)^2}{x^2}
\left(ux\right)'\:=1+u+u^2
Lest hand side \left(ux\right)'\:=xu'\:+u
xu'\:+u=1+u+u^2
\frac{1}{u^2+1}u'=\frac{1}{x}
\tan^{-1} \left(u)=\ln \left(x\right)+c_1
u=\tan \left(\ln \left(x\right)+c_1\right)
\mathrm{Substitute\:back}\:u=\frac{y}{x}
\frac{y}{x}=\tan \left(\ln \left(x\right)+c_1\right)